# Threshold Graphs are Globally Synchronizing

### v4

@@ Abstract.
The Kuramoto model describes phase oscillators on the unit circle whose interactions are encoded by a graph. Each edge acts like a spring that pulls adjacent oscillators toward each other whenever their phases differ. A central question is to determine which graphs are globally synchronizing, meaning that trajectories of the Kuramoto dynamics converge to the fully synchronized state from almost all initial conditions. This property is tightly linked to the benign nonconvexity of the model’s energy landscape. Existing guarantees for global synchronization typically require every node to connect to sufficiently many neighbors, which in turn enforces rather homogeneous degree sequences and highly dense graphs. In this work, we identify threshold graphs as a class of graphs that lie outside this regime yet nevertheless exhibit global synchronization. More precisely, connected threshold graphs realize all admissible edge densities ranging from $2/n$ to $1$, and their degree sequences are maximally left-skewed among all graphs with the same edge density, when ordered in nonincreasing order. Our proof relies on a phasor-geometric analysis of the stationary points of the associated energy landscape.
@@ 1. Introduction
The study of synchronization traces back to the 17th century, when Christiaan Huygens first observed the spontaneous synchronization of pendulum clocks. Since then, synchronization phenomena have been widely studied across both natural and technological systems. For example, the flashing of fireflies [BB68, SS93], the rhythmic beating of heart cells [Win67, MS06], synchronization in power grids [DB12], coordinated motion in robotic networks [OS06], modern artificial intelligence models such as neural networks [WCP+24], and Transformer architecture [GLPR25, CRMB24].
@@
$\frac{d\theta_{i}}{dt}=\omega_{i}+\sum_{j=1}^{n}A_{ij}\sin(\theta_{j}-\theta_{i}),\qquad i\in V,$ (1)
@@
Early studies of the Kuramoto model primarily considered systems of oscillators with heterogeneous intrinsic frequencies $\{\omega_{i}\}_{i=1}^{n}$, coupled through a complete graph with uniform weights $\boldsymbol{A}_{ij}=K/n$. Here, $K$ denotes the strength of the coupling between oscillators. A fundamental question in this setting is to determine the critical coupling strength $K_{c}$, such that for $K>K_{c}$, the oscillators converge to a common frequency, a phenomenon known as frequency synchronization [^1] That is, all derivatives $\tfrac{d\theta_{i}(t)}{dt}$ become identical as $t\to\infty$., and for $K<K_{c}$, they remain desynchronized. This problem is of central interest across many scientific domains, ranging from biological synchronization, rhythmic phenomena [MMJ87], technological applications such as deep brain stimulation [Tas03, NM11] and power systems [CCC95, DB12]. For comprehensive surveys of this line of work and its applications, we refer to [Str00, ABPV+05, DB14].
@@
$\frac{d\theta_{i}}{dt}=\sum_{j=1}^{n}A_{ij}\sin(\theta_{j}-\theta_{i}),\qquad i\in V.$ (2)
@@
Over the past two decades, a central question that has emerged is the following: under what graph structures does the homogeneous Kuramoto system achieve phase synchronization for all initial conditions, except for a set of measure zero? This problem is not only mathematically appealing but also closely connected to several recent developments across different research areas, including the Burer–Monteiro approach for semidefinite programming [BBV16, LXB19, Lin23] and the dynamics of Transformer architectures [GLPR25, GLPR23].
@@ 1.1. Problem Setting
We say that a graph $G=(V,E)$ is globally synchronizing if, for all initial conditions $\boldsymbol{\theta}(0)\in\mathbb{R}^{n}$ except for a set of measure zero, the solution to the Kuramoto model (2) on $G$ satisfies
@@
The key question is now stated as:
Which graphs are globally synchronizing?
@@
Historically, Verwoerd and Mason [VM07] conjectured in 2007 that every complete graph $K_{n}$ ($n\geq 2$) is globally synchronizing. This conjecture was resolved by Richard Taylor in 2012 [Tay12], and global synchronization of complete graphs has since become a standard benchmark result in this line of research. For classical results on the Kuramoto model on complete graphs, we refer to [ABPV+05, Str00, PR15, BGLM20].
In the same work [Tay12], Taylor also initiated what is now known as the extremal combinatorics version of the problem [BKMR25], which sparked substantial interest and subsequent developments. A more detailed discussion of these advances will be given in Prior Work.
One standard approach to this problem is to study the corresponding energy landscape. Note that the Kuramoto model (2) can be formulated as a gradient flow of a certain energy function. Precisely, the system of ordinary differential equations (2) defines a gradient dynamical system (or called gradient flow [^2] We call a dynamical system a gradient flow on $E(x)$ if it has the form $\frac{dx(t)}{dt}=-\nabla E(x)$. Its trajectories evolve along the steepest descent direction of $E$.) with energy function $E_{G}(\boldsymbol{\theta})$, satisfying
$\frac{d\boldsymbol{\theta}}{dt}=-\nabla E_{G}(\boldsymbol{\theta})$ (3)
@@
$=$ $\frac{1}{2}\sum_{1\leq i,j\leq n}\boldsymbol{A}_{ij}\big(1-\langle v_{i},v_{j}\rangle\big).$
Here, $v_{i}$ is the unit vector $(\cos\theta_{i},\sin\theta_{i})\in\mathbb{R}^{2}$.
Clearly, this function is nonconvex. Moreover, understanding the geometry of (4) provides direct insight into the long-term dynamics of (2). In particular, the synchronous states correspond exactly to the global minima of this energy function, and those minima that are local but not global (called spurious local minimua) correspond to non-synchronous stable equilibria. If such spurious local minima exist, the gradient dynamics may converge to them whenever the initialization lies within their basin of attraction, in which case the graph $G$ is not globally synchronizing. In fact, if one can show that the energy function (4) admits no spurious second-order stationary point (in such case we say the function is benign), then the graph $G$ is globally synchronizing. A rigorous proof of this statement can be established using the Łojasiewicz gradient inequality [Lag07] together with the center manifold theorem; see Lemma A.1 in [GLPR25] for an excellent exposition.
@@ 1.2. Prior Work
Previous works on globally synchronizing graphs primarily focus on identifying sufficient conditions under which a graph is guaranteed to exhibit global synchronization. A prominent line of research formulates this question in an extremal combinatorics version:
@@
The intuition for this problem is as follows. It is a classical result that complete graphs admit global synchronization [Tay12]. In this case, each node is connected to all other n1n-1 nodes, corresponding to the maximal value $\mu_{c}=1$. A natural question is whether the global synchronization property persists when $\mu_{c}=1$ is relaxed, that is, when each node is connected to only a fraction of the network. Several advances have been made in [Tay12, LXB19, LS21, KST21, Can22, YTT21]. The currently best known bounds on the critical minimum-degree threshold $\mu_{c}$ satisfy $0.6875\leq\mu_{c}\leq 0.75$, where the upper bound is established in [KST21], and the strongest known lower bound follows from [Can22]. It has been conjectured that the minimum-degree threshold for global synchronization is exactly $\mu_{c}=0.75$, and that below this density, even highly connected graphs may fail to synchronize (see [TSS20] and Conjecture 5 in [BKMR25]).
A remarkable feature of these minimum degree-based criteria is that, although they are only sufficient, they apply uniformly to all graphs meeting the prescribed threshold. In particular, global synchronization is guaranteed based solely on this parameter, without invoking any finer structural properties of the graph. Another important observation is that edge density alone is not an appropriate criterion for global synchronization. Indeed, Townsend, Stillman, and Strogatz [TSS20] provided examples of dense graphs that nevertheless fail to synchronize globally. Nevertheless, imposing a minimum degree condition necessarily forces the graph to be very dense. In particular, when $\mu_{c}\geq 75\%$, the total number of edges satisfies
$|E|\geq\frac{3}{8}\,n(n-1),$
which scales as $O(n^{2})$ and is therefore comparable to that of a complete graph.
At the opposite end of the connectivity spectrum from dense networks, how sparse a globally synchronizing graph can be?
Several existing works provide evidence toward this question. First, Townsend, Stillman, and Strogatz [TSS20] approached this problem by asking how few edges are required to transform a ring graph [^3] A ring graph of size $n\geq 5$ does not globally synchronize. into a globally synchronizing network. It was shown that destabilizing all twisted states, which form a family of non-synchronous equilibria of the ring graph on $n$ nodes, requires adding $O(n\log n)$ edges arranged in a specific manner. Numerical experiments further suggest that, for networks of modest size, no non-synchronous stable states persist after such modifications, indicating that certain sparse circulant graphs may exhibit global synchronization.
Second, results on globally synchronizing Erdős–Rényi graphs further indicate that networks substantially sparser than those satisfying the classical $75\%$ minimum-degree conditions can still exhibit global synchronization. Ling, Xu, and Bandeira [LXB19] first proved that global synchronization occurs with high probability as $n\to\infty$ provided that the edge probability satisfies $p\gg\log(n)/n^{1/3}$. This bound was later improved by Kassabov, Strogatz, and Townsend to $p\gg\log^{2}(n)/n$ [KST22]. Both works raised the conjecture that global synchronization should already occur as soon as the graph is connected, namely, for any $\varepsilon>0$ and $p\geq(1+\varepsilon)\frac{\log n}{n}$. This conjecture was recently resolved in [ABK+24], with further refinements obtained in [JMS25, McR25].
Taken together, the sparsest known families of globally synchronizing graphs beyond the trivial case of trees were supported by evidence at the $O(n\log n)$ edge scale. Whether global synchronization can be guaranteed for graphs with significantly fewer edges444beyond the trivial cases tree graphs remained mysterious.
@@ 1.3. Our Contribution
In this work, we identify a class of graphs that admit global synchronization without requiring the minimum degree to exceed prescribed thresholds.
@@
Figure 2. Four connected threshold graphs on 1919 vertices, from the sparsest ($S_{18}$) to the densest ($K_{19}$). Node labels indicate the construction order: vertex $1$ is the initial white vertex, and all subsequent labels follow the binary code of the construction. White and black circles denote isolated and dominating vertices (bit 0 and bit 11). Codes from left to right: 000000000000000001, 000100000000010001, 010001101101010001, $111111111$.
Notably, graphs in this class can have arbitrarily low or high edge density, ranging from $2/n$ to $1$, demonstrating that density alone is not an effective indicator of global synchronization. Moreover, their degree sequences are typically highly left-skewed, in sharp contrast to the relatively flat degree sequences arising in prior density-based and random graph models. This comparison further indicates that the shape of the degree sequence, by itself, is also not an appropriate indicator of global synchronization. We discuss the density and degree sequence of threshold graphs in Section 7.
In terms of proving the main result, the difficulty arises from the fact that many existing proofs of global synchronization rely, either explicitly or implicitly, on the so-called half-circle lemma. This lemma states that if all oscillators are confined to an open half-circle, then the dynamics contract toward synchrony. Building on this observation, the bootstrap strategy seeks to show that all second-order stationary points must lie within a half-circle. If successful, this excludes the existence of non-synchronous states to which the dynamics may converge, thereby implying global synchronization. However, for threshold graphs, it remains unclear how to apply this strategy. It typically depends on sufficient connectivity in the network thus naturally suited to dense graphs, where the large number of edges ensures that nodes inside the half-circle collectively exert sufficient force to pull outside nodes back into alignment. In contrast, threshold graphs span the full spectrum from extremely sparse examples such as stars to dense cliques, and their highly inhomogeneous, layered structure complicates the application of this strategy. As a result, existing techniques offer limited insight into whether and why these graphs synchronize globally.
@@
Understanding the Mechanism Behind Global Synchrony in Threshold Graphs. Previous global synchronization results focus on density-based conditions. For instance, any graph whose minimum degree exceeds 0.7889(n1)0.7889(n-1) is globally synchronizing [LS20]. In other words, these results require the graph to be sufficiently dense. Threshold graphs, however, lie outside this regime. Depending on the positions of dominating vertices in the generating sequence, their density can vary from $2/n$ (stars) to $1$ (cliques), and many intermediate densities occur. Nevertheless, we prove in this work that they also exhibit global synchronization.
@@ 1.5. Notation
Throughout this work, we consider only connected simple graphs, i.e., graphs without self-loops or weighted edges. Following the convention of dynamical systems, we refer to the variable $\theta=(\theta_{1},\cdots,\theta_{n})$ of the Kuramoto model (2) as a state, and call the solution $\boldsymbol{\theta}(t)$ with respect to the initial condition $\boldsymbol{\theta}(0)$ in the Kuramoto model (2) a trajectory starting from $\boldsymbol{\theta}(0)$. Moreover, $\mathbb{S}^{1}:=\{x\in\mathbb{R}^{2}:\|x\|=1\}$ denotes the unit circle. And $\boldsymbol{x}\Uparrow\boldsymbol{y}$ means that $\boldsymbol{x}$ and $\boldsymbol{y}$ are nonzero and positively collinear: $\boldsymbol{x}=\lambda\boldsymbol{y}$ for some $\lambda>0$. We write $u\sim v$ to indicate that vertices $u$ and $v$ are adjacent in a graph, and $u\nsim v$ otherwise. For a vertex $i$ and a subset $Q$ of vertices, we denote
$N_{Q}(i)=\{\,q\in Q:q\sim i\,\},$
the set of neighbors of $i$ that lie in $Q$.
@@ 2. Preliminaries of Threshold Graphs
Threshold graphs are a well-studied class of graphs that play a central role in graph theory and have found applications across a wide range of disciplines. They were discovered independently by researchers working in different fields and reported in separate venues. The term threshold graph was introduced by Chvátal and Hammer [CH73, CH77] in their study of set-packing problems. Around the same time, the same class of graphs was independently identified by Henderson and Zalcstein in the context of parallel processing, where they were referred to as PV-chunk definable graphs. Since then, threshold graphs have appeared in numerous settings, including resource allocation, cyclic scheduling, manpower planning, complexity theory, and the study of graphical degree sequences [MP95].
@@ 2.1. Definition and Characterizations
Although in this work we rely only on the construction process of threshold graphs to analyze their global synchronization, it is of independant interest to recall that they admit several equivalent characterizations [MP95].
@@
Interestingly, threshold graphs appeared in the synchronization literature long before their role in the Kuramoto model was considered. To the best of our knowledge, the earliest instance dates back to 1977, with the seminal work of Ausiello, Messina, and Protasi [AMP77], who provided a graph-theoretic characterization of the PVc class of synchronizing primitives and introduced the notion of PVc-definable graphs. This line of work emerged almost simultaneously with the independant introduction of threshold graphs by Chvátal and Hammer in their study of aggregation of inequalities in integer programming [CH77].
3. Preliminaries on Kuramoto Model
In this section, we introduce two families of concepts for the Kuramoto model: one arising from dynamical systems and the other from nonconvex optimization. Presenting these notions side by side is important, as many of the local synchronization lemmas developed later rely on interpreting stability to infer the relative positions of phasors at second-order stationary points of the energy landscape.
3.1. Equilibrium and Local Stability
In this section, we introduce the notion of equilibrium and stable equilibrium of the Kuramoto model (2).
By the definition, equilibrium points of a dynamical system are the points where the system remains at rest, meaning that the first-order derivative of the state vanishes [Kha02]. Setting the right-hand side of (2) to zero provides the characterization of an equilibrium point in the homogeneous Kuramoto model on the graph $G$.
Definition 3.1 (Equilibrium of the Kuramoto model (2)).

### v5

@@ Abstract.
The Kuramoto model can be formulated as a gradient flow on a nonconvex energy landscape of the form $E(\boldsymbol{\theta}):=\frac{1}{2}\sum_{1\leq i,j\leq n}A_{ij}\bigl(1-\cos(\theta_{i}-\theta_{j})\bigr).$ A fundamental question is to identify graph structures for which this landscape is benign, in the sense that every second-order stationary point corresponds to a fully synchronized state. This property guarantees that all trajectories of the Kuramoto model converge to a fully synchronized state except for a measure-zero set of initial conditions, a phenomenon known as global synchronization. Existing guarantees typically require that each node be connected to a sufficiently large fraction of the other nodes, enforcing high graph density. In this work, we show that threshold graphs lie well outside this regime while still exhibiting global synchronization. In particular, threshold graphs realize arbitrary edge densities and have degree sequences that are extremal in the sense of majorization. Our analysis is based on a phasor–geometric characterization of stationary points that exploits the structural and geometric symmetries induced by threshold graphs.
@@ 1. Introduction
The study of synchronization traces back to the 17th century, when Christiaan Huygens first observed the spontaneous synchronization of pendulum clocks. Since then, synchronization phenomena have been widely studied across both natural and technological systems. For example, the flashing of fireflies [BB68, SS93], the rhythmic beating of heart cells [Win67, MS06], synchronization in power grids [DB12], coordinated motion in robotic networks [OS06], modern artificial intelligence models such as neural networks [MLGW24], and Transformer architecture [GLPR25, CRMB24].
@@
$\frac{d\theta_{i}}{dt}=\omega_{i}+\sum_{j=1}^{n}A_{ij}\sin(\theta_{j}-\theta_{i}),\qquad i\in V,$
@@
Early studies of the Kuramoto model primarily considered systems of oscillators with heterogeneous intrinsic frequencies $\{\omega_{i}\}_{i=1}^{n}$, coupled through a complete graph with uniform weights $\boldsymbol{A}_{ij}=K/n$. Here, $K$ denotes the strength of the coupling between oscillators. A fundamental question in this setting is to determine the critical coupling strength $K_{c}$, such that for $K>K_{c}$, the oscillators converge to a common frequency, a phenomenon known as frequency synchronization [^1] That is, all derivatives $\tfrac{d\theta_{i}(t)}{dt}$ become identical as $t\to\infty$., and for $K<K_{c}$, they remain desynchronized. This problem is of central interest across many scientific domains, ranging from biological synchronization, rhythmic phenomena [MMJ87], technological applications such as deep brain stimulation [NM11] and power systems [CCC95, DB12]. For comprehensive surveys of this line of work and its applications, we refer to [Str00, ABPV+05, DB14].
@@
$\frac{d\theta_{i}}{dt}=\sum_{j=1}^{n}\boldsymbol{A}_{ij}\sin(\theta_{j}-\theta_{i}),\qquad i\in V.$ (1)
@@
Over the past two decades, a central question that has emerged is the following: under what graph structures does the homogeneous Kuramoto system achieve phase synchronization for all initial conditions, except for a set of measure zero? This problem is not only mathematically appealing but also closely connected to several recent developments across different research areas, including the Burer–Monteiro approach for semidefinite programming [BBV16, LXB19, Lin25] and the dynamics of Transformer architectures [GLPR25, GLPR23].
@@ 1.1. Problem Setting
We say that a graph $G=(V,E)$ is globally synchronizing if, for all initial conditions $\boldsymbol{\theta}(0)\in\mathbb{R}^{n}$, except for a set of measure zero, the solution to the Kuramoto model (1) on $G$ satisfies
@@
This naturally raises the following question:
Key question. Which graphs are globally synchronizing?
@@
1.2. Nonconvex Optimization Landscape
From an optimization viewpoint, global synchronization can be implied by the absence of spurious second-order stationary points of the nonconvex energy function. Precisely, the Kuramoto model (1) defines a gradient dynamical system (or called gradient flow [^2] We call a dynamical system a gradient flow on $E(x)$ if it has the form $\frac{dx(t)}{dt}=-\nabla E(x)$. Its trajectories evolve along the steepest descent direction of $E$.) with energy function $E_{G}(\boldsymbol{\theta})$, satisfying
$\frac{d\boldsymbol{\theta}}{dt}=-\nabla E_{G}(\boldsymbol{\theta})$
@@ 1.2. Nonconvex Optimization Landscape
$=$ $\frac{1}{2}\sum_{1\leq i,j\leq n}\boldsymbol{A}_{ij}\big(1-\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle\big).$
Here, $\mathbf{v}_{i}$ is the unit vector $(\cos\theta_{i},\sin\theta_{i})\in\mathbb{R}^{2}$.
Clearly, this function is nonconvex. Moreover, understanding the geometry of (2) provides direct insight into the long-term dynamics of (1). In particular, the synchronous states correspond exactly to the global minima of this energy function, and those minima that are local but not global (called spurious local minima) correspond to non-synchronous stable equilibria. If such spurious local minima exist, the gradient dynamics may converge to them whenever the initialization lies within their basin of attraction, in which case the graph $G$ is not globally synchronizing.
In contrast, if the energy function (2) admits no spurious second-order stationary points, in the sense that the only equilibria at which the Hessian $\nabla^{2}E_{G}(\boldsymbol{\theta})$ is positive semidefinite are the fully synchronized states, then the graph $G$ is globally synchronizing. A rigorous proof of this statement can be established using the Łojasiewicz gradient inequality [Lag07] together with the center manifold theorem; see Lemma A.1 in [GLPR25] for an excellent exposition.
@@ 1.3. Prior Work
Previous works on globally synchronizing graphs primarily focus on identifying sufficient conditions that guarantee global synchronization. A dominant line of research formulates this question from an extremal combinatorics perspective:
@@
Several advances have been made in [Tay12, LXB19, LS20, KST21, Can22, YTT21]. The currently best known bounds on the critical minimum-degree threshold $\mu_{c}$ satisfy $0.6875\leq\mu_{c}\leq 0.75$, where the upper bound is established in [KST21], and the strongest known lower bound follows from [Can22]. It has been conjectured that the minimum-degree threshold for global synchronization is exactly $\mu_{c}=0.75$, and that below this density, even highly connected graphs may fail to synchronize (see [TSS20] and Conjecture 5 in [BKMR25]).
Sufficiently large threshold for minimal degree enforces high density. Indeed, when $\mu_{c}\geq 75\%$, the total number of edges satisfies $|E|\geq\frac{3}{8}\,n(n-1)$, which scales as $O(n^{2})$. However, dense graphs do not necessarily global synchronize, as indicated by the examples in [TSS20]. How sparse a globally synchronizing graph can be? Townsend, Stillman, and Strogatz [TSS20] approached this problem by asking how few edges are required to transform a ring graph [^3] A ring graph of size $n\geq 5$ does not globally synchronize. into a globally synchronizing network. They showed that destabilizing all non-synchronizing twisted states requires adding $O(n\log n)$ edges arranged in a specific manner, indicating that adding $O(n\log n)$ edges to a ring graph in a strategic way may lead to global synchronization. Moreover, a line of work on global synchronization over random graphs [LXB19, KST22, ABK+26, JMS25, McR25] provides further evidence that sparse networks can globally synchronize. A natural question is to determine the critical connectivity threshold for Erdős–Rényi graphs to exhibit global synchronization with high probability. The best known bound asserts that, for any $\varepsilon>0$, global synchronization holds with high probability when $p\geq(1+\varepsilon)\frac{\log n}{n},$ as shown in [ABK+26], with further refinements in [JMS25, McR25]. However, this regime still corresponds to graphs with $O(n\log n)$ edges.
@@ 1.4. Our Contribution
In this work, we identify a class of graphs that admit global synchronization without requiring the minimum degree to exceed prescribed thresholds, and can have any edge density as will be shown in Section 7.
@@
Figure 2. Four connected threshold graphs on 1919 vertices, from the sparsest (star graph) to the densest (complete graph). Node labels indicate the vertex addition order in the construction, with vertex $1$ as the initial vertex. White and black circles denote isolated and dominating vertices (bit 0 and bit 11), and the initial vertex $1$ is colored in white. Codes from left to right: 000000000000000001, 000100000000010001, 010001101101010001, $111111111$.
In terms of proving the main result, the difficulty arises from the fact that many existing proofs of global synchronization rely, either explicitly or implicitly, on the so-called half-circle lemma. This lemma states that if all oscillators are confined to an open half-circle, then the dynamics contract toward synchrony. Building on this observation, the bootstrap strategy seeks to show that all second-order stationary points must lie within a half-circle. If successful, this excludes the existence of non-synchronous states to which the dynamics may converge, thereby implying global synchronization.
However, for threshold graphs, it remains unclear how to apply this strategy. It typically depends on sufficient connectivity in the network thus naturally suited to dense graphs, where the large number of edges ensures that nodes inside the half-circle collectively exert sufficient force to pull outside nodes back into alignment. In contrast, threshold graphs span the full spectrum from extremely sparse examples, such as stars, to extremely dense ones, such as complete graphs, making it unclear how such techniques can be applied in this setting.
@@
Understanding the Mechanism Behind Global Synchrony in Threshold Graphs. Previous global synchronization results focus on density-based conditions. For instance, any graph whose minimum degree exceeds 0.75(n1)0.75(n-1) is globally synchronizing [KST21]. In other words, these results require the graph to be sufficiently dense. Threshold graphs, however, lie outside this regime. Nevertheless, we prove in this work that they also exhibit global synchronization.
@@ 1.6. Notation
Throughout this work, all graphs are assumed to be connected and simple, i.e., without self-loops or weighted edges. We write $u\sim v$ to indicate that vertices $u$ and $v$ are adjacent in a graph, and $u\not\sim v$ otherwise. For a vertex $i$ and a subset $Q$ of vertices, we denote
$N_{Q}(i):=\{q\in Q:q\sim i\},$
[…]
We denote by $\mathbb{S}^{1}:=\{\boldsymbol{x}\in\mathbb{R}^{2}:\|\boldsymbol{x}\|=1\}$ the unit circle. For vectors $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{2}$, we write $\boldsymbol{x}\Uparrow\boldsymbol{y}$ to indicate that $\boldsymbol{x}$ and $\boldsymbol{y}$ are nonzero and positively collinear, i.e., $\boldsymbol{x}=\lambda\boldsymbol{y}$ for some $\lambda>0$.
Following the convention in dynamical systems, we refer to the variable $\boldsymbol{\theta}=(\theta_{1},\ldots,\theta_{n})$ of the Kuramoto model (1.2) as the state, and call the solution $\boldsymbol{\theta}(t)$ with initial condition $\boldsymbol{\theta}(0)$ a trajectory starting from $\boldsymbol{\theta}(0)$.
@@ 2. Preliminaries of Threshold Graphs
Threshold graphs are a well-studied class of graphs that play a central role in graph theory and have found applications across a wide range of disciplines. The term threshold graph was introduced by Chvátal and Hammer [CH77b] in their study of set-packing problems. Since then, threshold graphs have appeared in numerous settings, including resource allocation, cyclic scheduling, manpower planning, complexity theory, and the study of graphical degree sequences [CH77a].
@@ 2.1. Definition and Characterizations
Although in this work we rely only on the construction process of threshold graphs to analyze their global synchronization, it is of independent interest to recall that they admit several equivalent characterizations [CH77b].
@@
Interestingly, threshold graphs appeared in the synchronization literature long before their role in the Kuramoto model was considered. In particular, as early as 1977, Ausiello, Messina, and Protasi [AMP77] introduced a class of graphs (later recognized as threshold graphs) in the study of synchronization primitives in distributed computation. Their work provided a graph-theoretic characterization of so-called PVc-definable graphs, which encode combinatorial conditions under which simple local update rules lead to global agreement.
3. Preliminaries on the Kuramoto Model
3.1. Equilibria and Local Stability
By the definition, equilibrium points of a dynamical system are the states where the system remains at rest, meaning that the first-order derivative of the state vanishes [Kha02]. Thus setting the right-hand side of (1) to zero provides the characterization of an equilibrium point of the Kuramoto model.
Definition 3.1 (Equilibria of (1)).